OpenStudy (anonymous):

Solve: r5- 6 = -1 A) r = 1 B) r = 5 C) r = 10 D) r = 25

OpenStudy (mathmale):

Please explain (in words) what you mean by "r5." For example, are you referring to multiplication or to exponentiation? What is the goal of this problem?

OpenStudy (anonymous):

r/5

OpenStudy (mathmale):

Makes a big difference! So: r/5 - 6 = - 1. What step could you take to isolate r/5 on the left side?

OpenStudy (anonymous):

Yes thats right

OpenStudy (anonymous):

Add up the numbers

OpenStudy (mathmale):

Try adding 6 to both sides. Your result?

OpenStudy (anonymous):

1

OpenStudy (mathmale):

Where's the other side of your equation? Write out the whole equation, please.

OpenStudy (anonymous):

thats it

OpenStudy (mathmale):

Not so. You started with an equation and if your answer is correct, you'll end with an eqatuion. You began with r/5-6=-1 and added 6 to both sides. What does your equation look like now?

OpenStudy (anonymous):

I dont understand

OpenStudy (mathmale):

Let's write out your equation: r/5 - 6 = -1 Add 6 to both sides: +6 = +6 ____________________ r/5 = 5

OpenStudy (mathmale):

So now r/5 = 5. What would you do next, to isolate r?

OpenStudy (anonymous):

divide

OpenStudy (mathmale):

Multiply both sides of this latest equation by 5: \[5(r/5)=5(5)\]

OpenStudy (mathmale):

Please finish this work.

OpenStudy (mathmale):

Remember: multiplication is the inverse operation of division.

OpenStudy (mathmale):

Still interested in finishing this problem? If something's not clear to you, please ask questions as needed.

OpenStudy (anonymous):

Yea i dont understand...

OpenStudy (fedorable):

its A, r=1

KjSaif (kjsaif):

Hello no direct answers please :D

KjSaif (kjsaif):

First Add 6 to both sides r^5=−1+6

KjSaif (kjsaif):

Simplify −1+6−1+6 to 5 r^5=5

KjSaif (kjsaif):

Take the 5th root of both sides r=5√5

KjSaif (kjsaif):

which gives you 1.379 which rounded gets you 1 which gets you A :D

OpenStudy (mathmale):

We have a problem here: Original post was "r5-6=-1. I asked QueSmith for clarification. He said he meant r/5. Now you @KJSaif , are guessing that QueSmith meant 

KjSaif (kjsaif):

i don't know i said 5√5. Take the 5th root of both sides r=5√5 was i wrong?

OpenStudy (anonymous):

@KJSaif You told others to not give a direct answer, yet you gave a direct answer yourself right after that. Also, your answer was wrong. The asker (QueSmith) clarified that the question was: r/5 - 6 = -1 NOT r^5 - 6 = -1 And then you gave the answer first as \( 5\sqrt{5}\) when the correct answer to r^5 - 6 = -1 should actually be \(\sqrt[5]{5}\) Also, you were solving the wrong equation, so it really doesn't matter if the answer you gave was correct or wrong since it wouldn't answer the question given.

OpenStudy (anonymous):

Not to forget the fact that you gave an answer that wasn't even in the answer choices.

KjSaif (kjsaif):

i didn't give direct i explained

KjSaif (kjsaif):

maybe i messed up this question.

KjSaif (kjsaif):

stop

KjSaif (kjsaif):

wow thanks for ruining my day im leaving for a year because of YOU